# Beyond In-Domain Detection: SpikeScore for Cross-Domain Hallucination Detection

### v2

@@ Abstract
Hallucination detection is critical for deploying large language models (LLMs) in real-world applications. Existing hallucination detection methods achieve strong performance when the training and test data come from the same domain, but they suffer from poor cross-domain generalization. In this paper, we study an important yet overlooked problem, termed generalizable hallucination detection (GHD), which aims to train hallucination detectors on data from a single domain while ensuring robust performance across diverse related domains. In studying GHD, we simulate multi-turn dialogues following LLMs initial response and observe an interesting phenomenon: hallucination-initiated multi-turn dialogues universally exhibit larger uncertainty fluctuations than factual ones across different domains. Based on the phenomenon, we propose a new score SpikeScore, which quantifies abrupt fluctuations in multi-turn dialogues. Through both theoretical analysis and empirical validation, we demonstrate that SpikeScore achieves strong cross-domain separability between hallucinated and non-hallucinated responses. Experiments across multiple LLMs and benchmarks demonstrate that the SpikeScore-based detection method outperforms representative baselines in cross-domain generalization and surpasses advanced generalization-oriented methods, verifying the effectiveness of our method in cross-domain hallucination detection.
@@ 2 Learning Setups
Generalizable Hallucination Detection. In GHD, we consider a training domain $\mathbb{P}_{Q,T}$ and NN related test domains $\mathbb{P}_{Q,T}^{1},\ldots,\mathbb{P}_{Q,T}^{N}$. The goal of GHD is to train a detector on $\mathbb{P}_{Q,T}$ that generalizes well to the NN test domains. We formally introduce the definition of GHD below.
@@
as introduced in Eq. (2), the objective of generalizable hallucination detection (GHD) is to learn a detector $D$ based on the LLM $\mathbb{P}_{\bm{\theta}}(\cdot)$ and the training dataset $\mathcal{D}_{l}$. The detector is required to generalize across NN test domains $\mathbb{P}_{Q,T}^{1},\ldots,\mathbb{P}_{Q,T}^{N}$, i.e., for any question–answer pair $(\mathbf{Q},\mathbf{A})$,
@@ 3.1 Exploring Cross-Domain Invariant Indicators
Training Score for Uncertainity Estimation. To effectively evaluate the uncertainty of each continuation answer $\mathbf{A}^{k}$, we need to select suitable scoring methods. To further ensure that the scores remain relatively accurate in addressing hallucination detection, we adopt training-based methods that leverage information of training data tailored for hallucination detection.
In this work, we use the one of the most representative methods in hallucination detection, termed SAPLMA (azaria2023internal), which utilizes LLM’s internal state to reveal the truthfulness of a given answer. The details of SAPLMA is introduced as follows. Given the internal representation $\mathbf{E}_{\bm{\theta}}(\cdot)$ of the LLM $\mathbb{P}_{\bm{\theta}}(\cdot)$, we place a multilayer perceptron (MLP) followed by a sigmoid activation on top of $\mathbf{E}_{\bm{\theta}}(\cdot)$ to produce a probabilistic output $p_{\mathbf{W}}(\cdot)\in[0,1]$. The parameters $\mathbf{W}$ are then optimized using the cross-entropy loss over the training data $\mathcal{D}_{l}=\{(\mathbf{Q}_{i},\mathbf{A}_{i},y_{i})\}_{i=1}^{n}$, i.e.,
@@
where $K$ is the stoping step. In our experiment, we set $K=20$ as the defult.
@@
SpikeScore: Estimating Score Sequence Fluctuations. There are many methods to evaluate the sequence fluctuations, such as computing the variance or extreme gap. It is worth noting that sun2025detection proposes the coefficient of variation score, which measures the score sequence fluctuations using the coefficient of variation, i.e., the standard deviation normalized by the expection.
@@ 3.2 Evaluations of SpikeScore
Let $\mathbb{P}_{Q,T}^{t}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{P}_{Q,T}^{i}$ be the non-hallucination domain as the uniform mixture of NN test domains $\mathbb{P}_{Q,T}^{1},...,\mathbb{P}_{Q,T}^{N}$ introduced in Section 2. The value of SpikeScore trained on the associated training domain $\mathbb{P}_{Q,T}$ satisfies that
@@
In Figure 2 (b), we observe that the standard deviation of SpikeScore in the hallucination domain is larger than in the non-hallucination domain, with the ratio of standard deviations reaching $2.5$, which exceeds the expectation ratio of 22 reported in Observation 1. This indicates that variance differences alone may hinder direct separability. Nevertheless, as shown in Theorem 1, separability between hallucination and non-hallucination domains can still be established under controlled coefficient-of-variation conditions (i.e., variance normalized by expectation). See Theorem 1 for details.
@@ 3.3 Leveraging SpikeScore for GHD
In this part, we leverage SpikeScore to construct our method for GHD. Given a question $\mathbf{Q}$ and the corresponding answer $\mathbf{A}$, we first compute the SpikeScore of $\mathbf{A}$ according to Eq. (9). Based on this score, we then apply the following formulation to determine whether $\mathbf{A}$ is hallucinated or non-hallucinated: given a thershold $\lambda>0$,
@@
where 11 indicates that $\mathbf{A}$ is hallucinated, and 0 indicates that $\mathbf{A}$ is non-hallucinated.
@@ 4.1 Experimental Setups
We evaluate SpikeScore on six widely used hallucination detection benchmarks: CommonsenseQA (talmor-etal-2019-commonsenseqa), TriviaQA (joshi-etal-2017-triviaqa), Belebele (costa-jussa-etal-2025-2m), CoQA (reddy-etal-2019-coqa), Math (hendrycks2021measuring), and SVAMP (patel-etal-2021-nlp). Baselines include Perplexity (ren2023outofdistribution), Semantic Entropy (farquhar2024detecting), EigenScore (chen2024inside), Lexical Similarity (lin2024generating), Verbalize (lin2022teaching), InterrogateLLM (yehuda-etal-2024-interrogatellm), $M$ (marks2023geometry), SEP (kossen2024semantic; farquhar2024detecting), SAPLMA (azaria2023internal), PRISM (SAPLMA instantiation) (zhang2024prompt), and ICR Probe (zhang-etal-2025-icr), with AUROC as the primary evaluation metric. We test across four carefully selected open-source LLMs—Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct (abs-2407-21783), and Qwen3-8B-Instruct and Qwen3-14B-Instruct (qwen3)—to cover diverse model families and parameter scales. Additional dataset sampling strategies, baseline configurations, and hyperparameters are provided in detail in Appendix LABEL:app:details.
@@ 4.2 Experimental Results
Overall Cross-domain Hallucination Detection Performance. In Table 3.3, we compare the proposed SpikeScore with three categories of baselines: training-free methods that avoid generalization issues, traditional training-based methods, and cross-domain specialized approaches. SpikeScore consistently achieves the highest average AUROC across all model families and maintains superior or competitive performance in individual dataset evaluations. Traditional training-based methods ($M$, SEP, SAPLMA) generally exhibit poor cross-domain performance, while training-free methods show reasonable adaptability but fall short of specialized cross-domain approaches like PRISM and ICR Probe. Notably, SpikeScore demonstrates improved performance with larger models within the same family, suggesting that enhanced self-correction capabilities in larger LLMs (yang-etal-2025-confidence) generate more detectable attention spike patterns during conversational refinement.
@@ Appendix E Discussion on Separability
    Formally, for a single domain AA with hallucination and non-hallucination distributions $(X_{A},Y_{A})$, Cantelli’s inequality yields
@@
    where $\delta_{A}=\tfrac{\mathbb{E}X_{A}}{\mathbb{E}Y_{A}}$ is the mean ratio, $r_{A}=\tfrac{\mathrm{Std}(X_{A})}{\mathrm{Std}(Y_{A})}$ is the variance ratio, and $c_{A}=\tfrac{\mathrm{Std}(Y_{A})}{\mathbb{E}Y_{A}}$ is the coefficient of variation of non-hallucinations. If domain AA happens to exhibit weaker mean gaps ($\delta_{A}\!\approx\!1$), larger variance inflation ($r_{A}$ large), or higher non-hallucination noise ($c_{A}$ large), then its AUROC can drop below the global mixture bound. In practice, we indeed observe in Table 3.3 and Table E that datasets such as Belebele occasionally yield lower per-domain AUROCs relative to the theoretical bound. This does not contradict our analysis; rather, it reflects the fact that separability is guaranteed only at the mixture level, and domains with less favorable statistical profiles naturally appear weaker. Intuitively, these domains are “harder” because hallucinations and factual answers are less distinguishable in terms of score dynamics. Hence, the occasional $\mathrm{AUROC}_{A}<\mathrm{LB}_{\mathrm{mix}}$ is a theoretically anticipated phenomenon and does not necessitate further ablations.

### v3

@@ Abstract
Hallucination detection is critical for deploying large language models (LLMs) in real-world applications. Existing hallucination detection methods achieve strong performance when the training and test data come from the same domain, but they suffer from poor cross-domain generalization. In this paper, we study an important yet overlooked problem, termed generalizable hallucination detection (GHD), which aims to train hallucination detectors on data from a single domain while ensuring robust performance across diverse related domains. In studying GHD, we simulate multi-turn dialogues following LLMs’ initial response and observe an interesting phenomenon: hallucination-initiated multi-turn dialogues universally exhibit larger uncertainty fluctuations than factual ones across different domains. Based on the phenomenon, we propose a new score SpikeScore, which quantifies abrupt fluctuations in multi-turn dialogues. Through both theoretical analysis and empirical validation, we demonstrate that SpikeScore achieves strong cross-domain separability between hallucinated and non-hallucinated responses. Experiments across multiple LLMs and benchmarks demonstrate that the SpikeScore-based detection method outperforms representative baselines in cross-domain generalization and surpasses advanced generalization-oriented methods, verifying the effectiveness of our method in cross-domain hallucination detection.
@@ 2 Learning Setups
Generalizable Hallucination Detection. In GHD, we consider a training domain $\mathbb{P}_{Q,T}$ and $N$ related test domains $\mathbb{P}_{Q,T}^{1},\ldots,\mathbb{P}_{Q,T}^{N}$. The goal of GHD is to train a detector on $\mathbb{P}_{Q,T}$ that generalizes well to the $N$ test domains. We formally introduce the definition of GHD below.
@@
as introduced in Eq. (2), the objective of generalizable hallucination detection (GHD) is to learn a detector $D$ based on the LLM $\mathbb{P}_{\bm{\theta}}(\cdot)$ and the training dataset $\mathcal{D}_{l}$. The detector is required to generalize across $N$ test domains $\mathbb{P}_{Q,T}^{1},\ldots,\mathbb{P}_{Q,T}^{N}$, i.e., for any question–answer pair $(\mathbf{Q},\mathbf{A})$,
@@ 3.1 Exploring Cross-Domain Invariant Indicators
Training Score for Uncertainty Estimation. To effectively evaluate the uncertainty of each continuation answer $\mathbf{A}^{k}$, we need to select suitable scoring methods. To further ensure that the scores remain relatively accurate in addressing hallucination detection, we adopt training-based methods that leverage information of training data tailored for hallucination detection.
In this work, we use one of the most representative methods in hallucination detection, termed SAPLMA (azaria2023internal), which utilizes LLM’s internal state to reveal the truthfulness of a given answer. The details of SAPLMA is introduced as follows. Given the internal representation $\mathbf{E}_{\bm{\theta}}(\cdot)$ of the LLM $\mathbb{P}_{\bm{\theta}}(\cdot)$, we place a multilayer perceptron (MLP) followed by a sigmoid activation on top of $\mathbf{E}_{\bm{\theta}}(\cdot)$ to produce a probabilistic output $p_{\mathbf{W}}(\cdot)\in[0,1]$. The parameters $\mathbf{W}$ are then optimized using the cross-entropy loss over the training data $\mathcal{D}_{l}=\{(\mathbf{Q}_{i},\mathbf{A}_{i},y_{i})\}_{i=1}^{n}$, i.e.,
@@
where $K$ is the stopping step. In our experiment, we set $K=20$ as the default.
@@
SpikeScore: Estimating Score Sequence Fluctuations. There are many methods to evaluate the sequence fluctuations, such as computing the variance or extreme gap. It is worth noting that sun2025detection proposes the coefficient of variation score, which measures the score sequence fluctuations using the coefficient of variation, i.e., the standard deviation normalized by the expectation.
@@ 3.2 Evaluations of SpikeScore
Let $\mathbb{P}_{Q,T}^{t}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{P}_{Q,T}^{i}$ be the non-hallucination domain as the uniform mixture of $N$ test domains $\mathbb{P}_{Q,T}^{1},...,\mathbb{P}_{Q,T}^{N}$ introduced in Section 2. The value of SpikeScore trained on the associated training domain $\mathbb{P}_{Q,T}$ satisfies that
@@
In Figure 2 (b), we observe that the standard deviation of SpikeScore in the hallucination domain is larger than in the non-hallucination domain, with the ratio of standard deviations reaching $2.5$, which exceeds the expectation ratio of $2$ reported in Observation 1. This indicates that variance differences alone may hinder direct separability. Nevertheless, as shown in Theorem 1, separability between hallucination and non-hallucination domains can still be established under controlled coefficient-of-variation conditions (i.e., variance normalized by expectation). See Theorem 1 for details.
@@ 3.3 Leveraging SpikeScore for GHD
In this part, we leverage SpikeScore to construct our method for GHD. Given a question $\mathbf{Q}$ and the corresponding answer $\mathbf{A}$, we first compute the SpikeScore of $\mathbf{A}$ according to Eq. (9). Based on this score, we then apply the following formulation to determine whether $\mathbf{A}$ is hallucinated or non-hallucinated: given a threshold $\lambda>0$,
@@
where $1$ indicates that $\mathbf{A}$ is hallucinated, and 0 indicates that $\mathbf{A}$ is non-hallucinated.
@@ 4.1 Experimental Setups
We evaluate SpikeScore on six widely used hallucination detection benchmarks: CommonsenseQA (talmor-etal-2019-commonsenseqa), TriviaQA (joshi-etal-2017-triviaqa), Belebele (costa-jussa-etal-2025-2m), CoQA (reddy-etal-2019-coqa), Math (hendrycksmath2021), and SVAMP (patel-etal-2021-nlp). Baselines include Perplexity (ren2023outofdistribution), Semantic Entropy (farquhar2024detecting), EigenScore (chen2024inside), Lexical Similarity (lin2024generating), Verbalize (lin2022teaching), InterrogateLLM (yehuda-etal-2024-interrogatellm), MM (marks2023geometry), SEP (kossen2024semantic; farquhar2024detecting), SAPLMA (azaria2023internal), PRISM (SAPLMA instantiation) (zhang2024prompt), and ICR Probe (zhang-etal-2025-icr), with AUROC as the primary evaluation metric. We test across four carefully selected open-source LLMs—Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct (abs-2407-21783), and Qwen3-8B-Instruct and Qwen3-14B-Instruct (qwen3)—to cover diverse model families and parameter scales. Additional dataset sampling strategies, baseline configurations, and hyperparameters are provided in detail in Appendix LABEL:app:details.
@@ 4.2 Experimental Results
Overall Cross-domain Hallucination Detection Performance. In Table 3.3, we compare the proposed SpikeScore with three categories of baselines: training-free methods that avoid generalization issues, traditional training-based methods, and cross-domain specialized approaches. SpikeScore consistently achieves the highest average AUROC across all model families and maintains superior or competitive performance in individual dataset evaluations. Traditional training-based methods (MM, SEP, SAPLMA) generally exhibit poor cross-domain performance, while training-free methods show reasonable adaptability but fall short of specialized cross-domain approaches like PRISM and ICR Probe. Notably, SpikeScore demonstrates improved performance with larger models within the same family, suggesting that enhanced self-correction capabilities in larger LLMs (yang-etal-2025-confidence) generate more detectable attention spike patterns during conversational refinement.
@@ Appendix E Discussion on Separability
    Formally, for a single domain $A$ with hallucination and non-hallucination distributions $(X_{A},Y_{A})$, Cantelli’s inequality yields
@@
    where $\delta_{A}=\tfrac{\mathbb{E}X_{A}}{\mathbb{E}Y_{A}}$ is the mean ratio, $r_{A}=\tfrac{\mathrm{Std}(X_{A})}{\mathrm{Std}(Y_{A})}$ is the variance ratio, and $c_{A}=\tfrac{\mathrm{Std}(Y_{A})}{\mathbb{E}Y_{A}}$ is the coefficient of variation of non-hallucinations. If domain $A$ happens to exhibit weaker mean gaps ($\delta_{A}\!\approx\!1$), larger variance inflation ($r_{A}$ large), or higher non-hallucination noise ($c_{A}$ large), then its AUROC can drop below the global mixture bound. In practice, we indeed observe in Table 3.3 and Table E that datasets such as Belebele occasionally yield lower per-domain AUROCs relative to the theoretical bound. This does not contradict our analysis; rather, it reflects the fact that separability is guaranteed only at the mixture level, and domains with less favorable statistical profiles naturally appear weaker. Intuitively, these domains are “harder” because hallucinations and factual answers are less distinguishable in terms of score dynamics. Hence, the occasional $\mathrm{AUROC}_{A}<\mathrm{LB}_{\mathrm{mix}}$ is a theoretically anticipated phenomenon and does not necessitate further ablations.